OpenStudy (anonymous):

A residential lot has an area of 120 sq.m. If the length is 12 meters, what is its length?

OpenStudy (aravindg):

12!!

OpenStudy (anonymous):

AREA = Length * breadth

OpenStudy (aravindg):

are u sure the qn is right?

OpenStudy (aravindg):

do u need breadth?

OpenStudy (anonymous):

Great Job @Yahoo!

OpenStudy (anonymous):

120 = 12 * breadth breadth = 120/12 = 10 m

OpenStudy (aravindg):

breath =120/12=10

OpenStudy (anonymous):

Thanks. :D

OpenStudy (anonymous):

@BatmanBane abby?

OpenStudy (anonymous):

10